# Injecting Measurement Information Yields a Fast and Noise-Robust Diffusion-Based Inverse Problem Solver

### v1

@@ 4 Experiments
We run DCS and the other methods listed in Table 1 on the FFHQ-256 (Karras, Laine, and Aila 2019), (Kazemi and Sullivan 2014) and ImageNet (Deng et al. 2009) datasets. For the prior network $\epsilon_{\theta}$, we use the corresponding pretrained model weights from (Chung et al. 2022a).
@@
We find that DCS either outperforms, or is comparable to all existing methods. While some methods may excel at certain metrics in certain tasks and fail to recover the signal at other times, DCS consistently performs well across experiments. For example, DCS is one of few methods that has reasonable results on Motion Deblurring at high noise levels. DDNM and RED-Diff, on the other hand, are powerful across inpainting tasks in general, but fail to perform Motion Deblurring and show underwhelming qualitative performance on many tasks, especially in the high noise regime (Figures 5, 7). We find that projection methods in general overfit in the high noise regime. Other methods such as LGD-MC degrade less severely in the presence of noise, however do not perform as well in lower noise levels indicating underfitting across noise levels. It is apparant from both quantitative and qualitative comparisons that DCS and DPS strike a clearer balance between over- and underfitting.
We also notice that DCS provides a significant speedup and reduction in memory footprint compared to most methods, as noted in Table 1. We achieve this by not requiring backpropagation of the score network, as well as limiting the required number of neural function evaluations by using the measurement-consistent version of Tweedie’s formula.
@@ 4.1 Ablation on the Noise-aware Maximization Optimizer
We investigate how the choice of optimizer and parameters affects the noise-aware maximization algorithm in DCS. We note that the flexibility of using an optimizer enables us to make use of the stopping criterion as detailed in Section 3. In Figure 6 we run DCS with AdamW (Loshchilov, Hutter et al. 2017), SGD with momentum, and vanilla SGD to solve the SRx4 task on a subset of FFHQ. Runs of each optimizer at learning rate scaling factors are displayed to show the best performance, ensuring a fair comparison. It is clear in Figure 6 that the addition of a momentum term to the optimization process (both present in AdamW and SGD with momentum) can attain a higher level of image fidelity and solver stability than vanilla SGD. This provides empirical evidence for optimizer bias having an effect on solver performance in DCS. We see from this experiment that AdamW produces the most consistent results across learning rates, which motivates its use in our implementation.
@@ 4.2 Improvement on Jacobian-Free Implementations
A major gain in the empirical performance of DCS comes from the fact that it no longer requires backpropagations through the neural score function, which allows for reductions in both runtime and memory footprint. In theory, most diffusion-based solvers can be modified to remove this neural backpropagation step by applying a stop_gradient to the score function output (e.g., RED-Diff (Mardani et al. 2024)). We compare against RED-Diff in the main text, and additionally against ablated variants of DPS (Chung et al. 2022a) and LGD (Song et al. 2023b) in Appendix D.2, and demonstrate clear improvements on these methods.
@@ 5 Conclusion
We proposed an effective adjustment to the diffusion-based inverse problem solver framework in the literature that improves speed and stability. Observing that the marginals of the diffusion process which solves the inverse problem is Gaussian distributed at each time $t$, we derived a simple, single-parameter likelihood model, whose sole unknown variate is obtained via a tractable maximum likelihood estimation algorithm. We leveraged this new perspective to create a noise-aware maximization scheme, and demonstrated the effectiveness of our method via a suite of numerical experiments and qualitative comparisons.
@@ A.1 Diffusion Models
Inspired by non-equilibrium thermodynamics, denoising diffusion probabilistic models (Ho, Jain, and Abbeel 2020) convert data $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$ to noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ via a diffusion process described by the variance-preserving stochastic differential equation (VP-SDE)
@@
where $\beta(t):\mathbb{R}\rightarrow[0,1]$ is a monotonically increasing noise schedule and $\mathbf{w}$ is the standard Wiener process (Song et al. 2020). This leads to the marginal distribution
@@
where $\mathcal{N}(\ \cdot\ ;\mu,\Sigma)$ is the probability density function (pdf) of a normal distribution centered at $\mu$ with covariance $\Sigma$. Sampling from $p_{\text{data}}(\mathbf{x})$ can then occur by modeling the reverse diffusion, which has a simple form given by (Anderson 1982)
@@
and sampling consists of obtaining solutions to the reverse-time SDE (17) with numerical solvers. A simple approach is given by the DDIM sampler with $\sigma_{t}=\sqrt{1-\alpha_{t}}$ (Song, Meng, and Ermon 2020)
@@ A.2 Solving Inverse Problems with Diffusion Models
and describes the core method of the DPS algorithm (Chung et al. 2022a). This strategy can also be extended to latent diffusion models, resulting in Latent-DPS and PSLD (Rout et al. 2023). Generally, the conditional term $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$ cannot be exact due to reasons we will investigate subsequently in Section 2, though these approximations are improved in LGD (Song et al. 2023b) and STSL (Rout et al. 2024). More recent work (Sun et al. 2024) propose an annealed Monte-Carlo-based perspective to posterior sampling, which results in a very similar algorithm to DPS. Much like MCG and ReSample (discussed in the next category), posterior solvers require estimating $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagation through the diffusion model, and significantly increases runtime and hampers scalability compared to unconditional sampling.
@@
Where $\hat{\mathbf{x}}_{0}[\mathbf{x}_{t}]$ is some prediction of $\mathbf{x}_{0}$ given only $\mathbf{x}_{t}$ (we elaborate in Section 2), and $\mathbf{P}$ is either a projection onto the low rank subspace or range of $\mathcal{A}$. The resulting algorithms are DDRM (Kawar et al. 2022) and DDNM (Wang, Yu, and Zhang 2022), respectively. Of course, this strategy is often restricted to situations where two conditions simultaneously hold true: (1) the measurement operator $\mathcal{A}$ is linear, and (2) the inverse problem is noiseless, i.e, $\eta$ is identically $\mathbf{0}$. These assumptions drastically limit the applicability of such models. The linearity restriction can be lifted by taking derivatives the measurement discrepancy, as in MCG (Chung et al. 2022b) and ReSample (Song et al. 2024), though this comes at the cost of significantly increased computation, requiring $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagating through the score network. Finally, (Cardoso et al. 2023) straddles the line between both categories — while MCGdiff is ostensibly a Bayesian solver, it bears greater resemblance to projection solvers since it does not form the decomposition in Eq. 22 and also samples by projecting each iterate to the null-space of $\mathcal{A}$, thus implementing a projected n-particle sequential monte carlo (SMC) sampling algorithm.
@@ A Maximum Likelihood Solver
We take a different perspective on solving the inverse problem. As seen in Section 2, both projection and posterior solvers must quantify the discrepancy between $\mathbf{x}_{t}$ and $\mathbf{y}$ via the consistency error $\mathcal{A}(\mathbf{x}_{0})-\mathbf{y}||$ at each diffusion step. Due to the complexity of the diffusion process, this involves approximating a fundamentally intractable quantity. In Section 3, we construct a simpler process whose parameters can be obtained via maximum likelihood estimation. Unlike the evidence lower bound proposed in (Mardani et al. 2024), we derive an explicit likelihood model, which is amenable to an optimization scheme with a probabilistic noise-aware stopping criterion. Finally, we show that the resulting algorithm is simple, fast, and adaptable to noise.
@@ B.1 Proof of Tweedie’s Formula
$p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})=\frac{1}{\alpha_{t}}p\left(\alpha_{t}^{-1}\mathbf{x}\right)$ (28)
@@
$\frac{\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t},\mathbf{y}]-\mathbf{x}_{t}}{\sigma_{t}^{2}}$ $=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}p(\mathbf{x}^{\prime}|\mathbf{x}_{t})d\mathbf{x}^{\prime}$
@@
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{1}{\alpha_{t}}\left[\mathbf{x}_{t}+\sigma_{t}^{2}\nabla\log p_{t}(\mathbf{x}_{t}|\mathbf{y})\right].$ (30)
@@
First, we note that Tweedie’s formula (Efron 2011) tells us that the posterior mean of a data distribution $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ can be obtained via the relation
@@
$=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{p(\mathbf{x}^{\prime},\mathbf{x}_{t}|\mathbf{y})}{p(\mathbf{x}_{t}|\mathbf{y})}d\mathbf{x}^{\prime}$
$=\int\frac{\alpha_{t}\mathbf{x}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})d\mathbf{x}}d\mathbf{x}^{\prime}$
@@ D.1 Comparison against Latent Models (Table 4)
We show that our pixel-based model also performs favorably against latent models in Table 4. We retain the same experimental setting on pixel-based models as in Table 2. For FFHQ, we use the pretrained FFHQ model weights from (Chung et al. 2022a) for our method, and the pretrained FFHQ model with a VQ-F4 first stage model (Rombach et al. 2022) in latent space models. For ImageNet, we again use pretrained model weights from (Chung et al. 2022a) in pixel-based diffusion solvers, and the Stable Diffusion v1.5 latent model for latent solvers. As with pixel-based methods many existing works suffer in the presence of additional noise. Further implementation details are discussed in Appendix E.
@@
Latent-DPS (Chung et al. 2023)  [^3] Latent-DPS is a direct application of DPS (Chung et al. 2023) to latent diffusion models. It is also mentioned in (Rout et al. 2023). Posterior Latent $6.1\mathbf{x}$ $8.9\mathbf{x}$
PSLD (Rout et al. 2023) Posterior Latent $7.5\mathbf{x}$ $15\mathbf{x}$
STSL (Rout et al. 2024) Posterior Latent $1.85\mathbf{x}$ $9\mathbf{x}$
ReSample (Song et al. 2024) Projection Latent  [^4] As described in (Song et al. 2024), ReSample does not run backpropagation on the score network, however the implementation does (Appendix E.4). $29.5\mathbf{x}$ $8.95\mathbf{x}$
DPS-JF (Chung et al. 2023) Posterior Pixel $5\mathbf{x} 1.1$$1\mathbf{x}$
LGD-MC (n=10) (Song et al. 2023b) Posterior Pixel 6x 3.2x
LGD-MC-JF (n=10) (Song et al. 2023b) Posterior Pixel $2\mathbf{x}$ $1.1\mathbf{x}$

### v2

@@ 4 Experiments
We run DCS and the other methods listed in Table 1 on the FFHQ-256 (Karras et al.,, 2019), (Kazemi and Sullivan,, 2014) and ImageNet (Deng et al.,, 2009) datasets. For the prior network $\epsilon_{\theta}$, we use the corresponding pretrained model weights from (Chung et al., 2022a, ).
@@
We find that DCS either outperforms, or is comparable to all existing methods. While some methods may excel at certain metrics in certain tasks and fail to recover the signal at other times, DCS consistently performs well across experiments. For example, DCS is one of few methods that has reasonable results on Motion Deblurring at high noise levels. DDNM and RED-Diff, on the other hand, are powerful across inpainting tasks in general, but fail to perform Motion Deblurring and show underwhelming qualitative performance on many tasks, especially in the high noise regime (Figures 5, 7). We find that projection methods in general overfit in the high noise regime. Other methods such as LGD-MC degrade less severely in the presence of noise, however do not perform as well in lower noise levels indicating underfitting across noise levels. It is apparant from both quantitative and qualitative comparisons that DCS and DPS strike a clearer balance between over- and underfitting. We also notice that DCS provides a significant speedup and reduction in memory footprint compared to most methods, as noted in Table 1. We achieve this by not requiring backpropagation of the score network, as well as limiting the required number of neural function evaluations by using the measurement-consistent version of Tweedie’s formula.
@@ 4.1 Ablation on the Noise-aware Maximization Optimizer
We investigate how the choice of optimizer and parameters affects the noise-aware maximization algorithm in DCS. We note that the flexibility of using an optimizer enables us to make use of the stopping criterion as detailed in Section 3. In Figure 6 we run DCS with AdamW Loshchilov et al., (2017), SGD with momentum, and vanilla SGD to solve the SRx4 task on a subset of FFHQ. Runs of each optimizer at learning rate scaling factors are displayed to show the best performance, ensuring a fair comparison. It is clear in Figure 6 that the addition of a momentum term to the optimization process (both present in AdamW and SGD with momentum) can attain a higher level of image fidelity and solver stability than vanilla SGD. This provides empirical evidence for optimizer bias having an effect on solver performance in DCS. We see from this experiment that AdamW produces the most consistent results across learning rates, which motivates its use in our implementation.
@@ 4.2 Improvement on Jacobian-Free Implementations
A major gain in the empirical performance of DCS comes from the fact that it no longer requires backpropagations through the neural score function, which allows for reductions in both runtime and memory footprint. In theory, most diffusion-based solvers can be modified to remove this neural backpropagation step by applying a stop_gradient to the score function output (e.g., RED-Diff (Mardani et al.,, 2024)). We compare against RED-Diff in the main text, and additionally against ablated variants of DPS (Chung et al., 2022a, ) and LGD (Song et al., 2023b, ) in Appendix D.2, and demonstrate clear improvements on these methods.
5 Impact and Limitations
Our work proposes a substantial improvement to the framework of noise-robust guided generation via diffusion models. As shown in Section 4 and Appendix D, DCS achieves consistent gains in performance and robustness, while also reducing runtime and memory requirements. These advances, along with the availability of our code, make the method particularly useful for researchers and practitioners with limited computational budgets, positioning DCS as a practical and readily adoptable tool for users working with inverse problems.
Our experiments focus on inverse problems with linear corrupting operators $\mathcal{A}$. While our theoretical framework extends to a broader class of non-linear operators (Appendix B.4), further investigation is required to fully understand the behavior of DCS and competing methods in non-linear settings.
@@ 6 Conclusion
We proposed an effective adjustment to the diffusion-based inverse problem solver framework in the literature that improves speed and stability. Observing that the marginals of the diffusion process, which solve the inverse problem, are normally distributed at each time $t$, we derived a simple, single-parameter likelihood model, whose sole unknown is obtained via a tractable maximum likelihood estimation algorithm. We leveraged this new perspective to create a noise-aware maximization scheme, and demonstrated the effectiveness of our method with a suite of numerical experiments and qualitative comparisons.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes, see link in footnote]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes, proofs are in the supplementary document.]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Injecting Measurement Information Yields a Fast and Noise-Robust Diffusion-Based Inverse Problem Solver:
Supplementary Materials
@@ A.1 Diffusion Models
Inspired by non-equilibrium thermodynamics, denoising diffusion probabilistic models (Ho et al.,, 2020) convert data $\mathbf{x}_{0}\sim p_{\text{data}}(\mathbf{x})$ to noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ via a diffusion process described by the variance-preserving stochastic differential equation (VP-SDE)
@@
where $\beta(t):\mathbb{R}\rightarrow[0,1]$ is a monotonically increasing noise schedule and $\mathbf{w}$ is the standard Wiener process (Song et al., 2020b, ). This leads to the marginal distribution
@@
where $\mathcal{N}(\ \cdot\ ;\mu,\Sigma)$ is the probability density function (pdf) of a normal distribution centered at $\mu$ with covariance $\Sigma$. Sampling from $p_{\text{data}}(\mathbf{x})$ can then occur by modeling the reverse diffusion, which has a simple form given by (Anderson,, 1982)
@@
and sampling consists of obtaining solutions to the reverse-time SDE (17) with numerical solvers. A simple approach is given by the DDIM sampler with $\sigma_{t}=\sqrt{1-\alpha_{t}}$ (Song et al., 2020a, )
@@ A.2 Solving Inverse Problems with Diffusion Models
and describes the core method of the DPS algorithm (Chung et al., 2022a, ). This strategy can also be extended to latent diffusion models, resulting in Latent-DPS and PSLD (Rout et al.,, 2023). Generally, the conditional term $\nabla\log p(\mathbf{y}|\mathbf{x}_{t})$ cannot be exact due to reasons we will investigate subsequently in Section 2, though these approximations are improved in LGD (Song et al., 2023b, ) and STSL (Rout et al.,, 2024). More recent work (Sun et al.,, 2024) propose an annealed Monte-Carlo-based perspective to posterior sampling, which results in a very similar algorithm to DPS. Much like MCG and ReSample (discussed in the next category), posterior solvers require estimating $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagation through the diffusion model, and significantly increases runtime and hampers scalability compared to unconditional sampling.
@@
Where $\hat{\mathbf{x}}_{0}[\mathbf{x}_{t}]$ is some prediction of $\mathbf{x}_{0}$ given only $\mathbf{x}_{t}$ (we elaborate in Section 2), and $\mathbf{P}$ is either a projection onto the low rank subspace or range of $\mathcal{A}$. The resulting algorithms are DDRM (Kawar et al.,, 2022) and DDNM (Wang et al.,, 2022), respectively. Of course, this strategy is often restricted to situations where two conditions simultaneously hold true: (1) the measurement operator $\mathcal{A}$ is linear, and (2) the inverse problem is noiseless, i.e, $\eta$ is identically $\mathbf{0}$. These assumptions drastically limit the applicability of such models. The linearity restriction can be lifted by taking derivatives the measurement discrepancy, as in MCG (Chung et al., 2022b, ) and ReSample (Song et al.,, 2024), though this comes at the cost of significantly increased computation, requiring $\frac{\partial}{\partial\mathbf{x}_{t}}\mathbf{x}_{0}$ which involves backpropagating through the score network. Finally, (Cardoso et al.,, 2023) straddles the line between both categories — while MCGdiff is ostensibly a Bayesian solver, it bears greater resemblance to projection solvers since it does not form the decomposition in Eq. 22 and also samples by projecting each iterate to the null-space of $\mathcal{A}$, thus implementing a projected n-particle sequential monte carlo (SMC) sampling algorithm.
@@ A Maximum Likelihood Solver
We take a different perspective on solving the inverse problem. As seen in Section 2, both projection and posterior solvers must quantify the discrepancy between $\mathbf{x}_{t}$ and $\mathbf{y}$ via the consistency error $\mathcal{A}(\mathbf{x}_{0})-\mathbf{y}||$ at each diffusion step. Due to the complexity of the diffusion process, this involves approximating a fundamentally intractable quantity. In Section 3, we construct a simpler process whose parameters can be obtained via maximum likelihood estimation. Unlike the evidence lower bound proposed in (Mardani et al.,, 2024), we derive an explicit likelihood model, which is amenable to an optimization scheme with a probabilistic noise-aware stopping criterion. Finally, we show that the resulting algorithm is simple, fast, and adaptable to noise.
@@ B.1 Proof of Tweedie’s Formula
$p_{\alpha_{t}}(\mathbf{x})=\frac{1}{\alpha_{t}}p\left(\alpha_{t}^{-1}\mathbf{x}\right)$ (28)
@@
$\frac{\mathbb{E}[\mathbf{x}^{\prime}|\mathbf{x}_{t}]-\mathbf{x}_{t}}{\sigma_{t}^{2}}$ $=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}p(\mathbf{x}^{\prime}|\mathbf{x}_{t})d\mathbf{x}^{\prime}$
@@
First, we note that Tweedie’s formula (Efron,, 2011) tells us that the posterior mean of a data distribution $\mathbf{x}_{t}\sim p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ can be obtained via the relation
@@
$=\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{p(\mathbf{x}^{\prime},\mathbf{x}_{t}|\mathbf{y})}{p(\mathbf{x}_{t}|\mathbf{y})}d\mathbf{x}^{\prime}.$
We now substitute $p(\mathbf{x}^{\prime},\mathbf{x}_{t}|\mathbf{y})=p(\mathbf{x}^{\prime}-\mathbf{x}_{t}|\mathbf{y})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})=\phi_{\sigma_{t}}(\mathbf{x}^{\prime}-\mathbf{x}_{t})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})$, due to the fact that the quantity $\mathbf{x}^{\prime}-\mathbf{x}_{t}\sim\mathcal{N}(0,\sigma\mathbf{I})$ is independent of $\mathbf{x}^{\prime}$ and $\mathbf{y}$. Hence, we continue the derivation
$\int\frac{\mathbf{x}^{\prime}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{p(\mathbf{x}^{\prime},\mathbf{x}_{t}|\mathbf{y})}{p(\mathbf{x}_{t}|\mathbf{y})}d\mathbf{x}^{\prime}$ $=\int\frac{\alpha_{t}\mathbf{x}_{0}-\mathbf{x}_{t}}{\sigma_{t}^{2}}\frac{\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x}^{\prime})p_{\alpha_{t}}(\mathbf{x}^{\prime}|\mathbf{y})}{\int\phi_{\sigma_{t}}(\mathbf{x}_{t}-\mathbf{x})p_{\alpha_{t}}(\mathbf{x}|\mathbf{y})d\mathbf{x}}d\mathbf{x}^{\prime}$
@@ D.1 Comparison against Latent Models (Table 4)
We show that our pixel-based model also performs favorably against latent models in Table 4. We retain the same experimental setting on pixel-based models as in Table 2. For FFHQ, we use the pretrained FFHQ model weights from (Chung et al., 2022a, ) for our method, and the pretrained FFHQ model with a VQ-F4 first stage model (Rombach et al.,, 2022) in latent space models. For ImageNet, we again use pretrained model weights from (Chung et al., 2022a, ) in pixel-based diffusion solvers, and the Stable Diffusion v1.5 latent model for latent solvers. As with pixel-based methods many existing works suffer in the presence of additional noise. Further implementation details are discussed in Appendix E.
@@
Latent-DPS (Chung et al.,, 2023)  [^3] Latent-DPS is a direct application of DPS Chung et al., (2023) to latent diffusion models. It is also mentioned in (Rout et al.,, 2023). Posterior Latent $6.1\mathbf{x}$ $8.9\mathbf{x}$
PSLD (Rout et al.,, 2023) Posterior Latent $7.5\mathbf{x}$ $15\mathbf{x}$
STSL (Rout et al.,, 2024) Posterior Latent $1.85\mathbf{x}$ $9\mathbf{x}$
ReSample (Song et al.,, 2024) Projection Latent  [^4] As described in (Song et al.,, 2024), ReSample does not run backpropagation on the score network, however the implementation does (Appendix E.4). $29.5\mathbf{x}$ $8.95\mathbf{x}$
DPS-JF (Chung et al.,, 2023) Posterior Pixel $5\mathbf{x} 1.1$$1\mathbf{x}$
LGD-MC (n=10) Song et al., 2023b Posterior Pixel 6x 3.2x
LGD-MC-JF (n=10) Song et al., 2023b Posterior Pixel $2\mathbf{x}$ $1.1\mathbf{x}$